# Cutoff for Contingency Table Random Walks

### v1

@@ 6.3 Bounding Decay in Characteristic Function for Small $\theta$
In this section, we bound the decay of the characteristic function $\Phi_{Y_{1}}(\theta)$ in the regime $\|\theta\|_{\infty}\leq\dfrac{1}{\sqrt{q}}$.
@@ 6.3.1 The Schur Complement of a Submatrix of $\Psi$
For convenience, we write $A:=\dfrac{n^{2}}{(n-1)^{2}}\Psi$ and thus
@@
Fix any $\ell,k\in\{1,\dots,n-1\}$. Let $A^{*}=A^{*}(\ell,k)$ be the principal submatrix of $A$ that only contains indices from the first $\ell$ rows and $k$ columns of the contingency table. We further partition $A^{*}$ as
@@
Let $A^{*}=A^{*}(\ell,k)$ be as above. Then the Schur complement $A^{*}_{22}-A^{*}_{21}(A^{*}_{11})^{-1}A^{*}_{12}$ satisfies
@@
We break the calculation into three cases. First, we consider an index $(i,k)$, $i\neq\ell$. As indicated in Figure 3, within the first $\ell$ rows and $k$ columns of the contingency table, there exist $\ell-2$ indices in the $k$th column and a row not the $i$th or the $\ell$th, $k-1$ indices in the $i$th row and a column not the $k$th, $k-1$ indices in the $\ell$th row and a column not the $k$th, and $(k-1)(\ell-2)$ indices that are neither in the $i$th or the $\ell$th row, nor in the $k$th column. A direct calculation then yields
@@
Figure 3: Diagram for an index $(i,k)$ with a shared $k$th column with $(\ell,k)$ in the submatrix containing the first $\ell$ rows and $k$ columns of the $n\times n$ contingency table. Note that this diagram is only for counting purposes, and the rows and columns may have been permuted.
Next, we deal with the case with an index $(\ell,j)$ with $j\neq k$. Analogously, in the first $\ell$ rows and $k$ columns of the contingency table, there are $k-2$ indices in the $\ell$th row and a column not the $j$th or the $k$th, $\ell-1$ indices in the $k$th column and a row not the $\ell$th, $\ell-1$ indices in the $j$th column and a row not the $\ell$th, and $(\ell-1)(k-2)$ indices neither in the $\ell$th row nor in the $k$th or the $j$th column. Therefore,
@@
In the last case, we check for an index $(i,j)$ with $i\neq\ell$, $j\neq k$. As in Figure 4, in the first $\ell$ rows and $k$ columns of the contingency table, there are $\ell-2$ indices in the $k$th column and a row not the $\ell$th or the $i$th, $k-2$ indices in the $\ell$th row and a column not the $k$th or the $j$th, $\ell-2$ indices in the $j$th column and a row not the $\ell$th or the $i$th, $k-2$ indices in the $i$th row and a column not the $k$th or the $j$th, and $(\ell-2)(k-2)$ indices that are neither in the $\ell$th or $i$th row nor in the $k$th or the $j$th column. This then gives
$(A^{*}_{11}v)_{(i,j)}=-\frac{4}{\ell k}+\frac{2}{\ell}+\frac{2}{k}+\frac{\ell-2}{\ell}+\frac{k-2}{k}-\frac{2(\ell-2)}{\ell k}-\frac{2(k-2)}{\ell k}-\frac{(\ell-2)(k-2)}{\ell k}=1.$
@@
Figure 4: Diagram for an index $(i,j)$ that does not share a row or a column with $(\ell,k)$ in the submatrix containing the first $\ell$ rows and $k$ columns of the $n\times n$ contingency table. Note that this diagram is only for counting purposes, and the rows and columns may have been permuted.
Therefore, we have $A^{*}_{11}v=A^{*}_{12}$ and thus $(A^{*}_{11})^{-1}A^{*}_{12}=v$. Now since there are $k-1$ indices with row $\ell$ other than $(\ell,k)$, and $\ell-1$ indices with column $k$ other than $(\ell,k)$, a direct computation yields
@@ 6.3.2 The Correlation Condition
We now show that the correlation matrices $\Gamma_{n}$ for the random walk on the $n\times n$ contingency table meet the Correlation Condition.
@@
Let $(\Gamma_{n})_{n=1}^{\infty}$ be the sequence of correlation matrices corresponding to the contingency table walk on an $n\times n$ table. Then $\Gamma_{n}$ meets the Correlation Condition with bounding function
$g(\alpha)=\frac{1+2e^{4/e}+2e^{8/e}+e^{32/e}}{\alpha^{1/4}}.$
@@
$\Psi_{n}((i,j),(i^{\prime},j^{\prime}))=\begin{cases}\frac{4(n-1)^{2}}{n^{2}}&(i,j)=(i^{\prime},j^{\prime})\\ \frac{2(n-1)^{2}}{n^{2}}&i=i^{\prime}\text{ or }j=j^{\prime}\ \text{but not both}\\ \frac{(n-1)^{2}}{n^{2}}&i\neq i^{\prime}\text{ and }j\neq j^{\prime}\end{cases}.$
It follows that $\sup_{n}\|\Psi_{n}\|_{\infty}=4$ which confirms condition (i). In order to check condition (ii), we first specify an ordering of the elements of $\Psi_{n}$ in order to compute each respective Schur complement: the first $n-1$ elements will consist of the first row of the table (incrementing by column index), second $n-1$ elements will consist of row two in the same order, continuing for each subsequent row.
@@
such that $M_{n,(\ell,k)}^{(22)}\in{\mathbb{R}}$ corresponds to element $(\ell,k)$. Then, since $M_{n,(\ell,k)}$ is a principal submatrix of $\Psi_{n,(\ell,k)}$, by a similar decomposition as (35),
@@
$M_{n,(\ell,k)}^{(22)}-M_{n,(\ell,k)}^{(21)}\left(M_{n,(\ell,k)}^{(11)}\right)^{-1}M_{n,(\ell,k)}^{(12)}=\left(\frac{n-1}{n}\right)^{2}\left(1+\frac{\ell+k+1}{\ell k}\right).$
@@
Now if either $\ell=1$ or $k=1$, say $\ell=1$, then note that $\Psi_{n,(1,k)}=\dfrac{2(n-1)^{2}}{n^{2}}(I_{k}+J_{k})$. So by (39),
$a_{\ell,k}^{(n)}=\left(\Psi_{n,(\ell,k)}^{(22)}-\Psi_{n,(\ell,k)}^{(21)}\left(\Psi_{n,(\ell,k)}^{(11)}\right)^{-1}\Psi_{n,(\ell,k)}^{(12)}\right)^{-1}=\frac{\ell k}{(\ell+1)(k+1)}$
@@
$\sum_{\ell=1}^{n}\sum_{k=1}^{n}\left(\frac{1}{\alpha n^{2}}\right)^{a_{\ell,k}^{(n)}}$ $\leq\sum_{\ell=1}^{n}\sum_{k=1}^{n}\left(\frac{1}{\alpha n^{2}}\right)^{\frac{\ell k}{(\ell+1)(k+1)}}$
$\leq\frac{1}{\alpha^{1/4}}\sum_{\ell=1}^{n}\sum_{k=1}^{n}\left(\frac{1}{n^{2}}\right)^{\frac{\ell k}{(\ell+1)(k+1)}}$
$\leq\frac{1+2e^{4/e}+2e^{8/e}+e^{32/e}}{\alpha^{1/4}}$
@@
$t=\dfrac{n^{2}q^{2}\log(\sqrt{\alpha}(n-1))}{8\pi^{2}}\left(1-\dfrac{4}{3q}\right)^{-1}.$
@@
$\sum_{y\in{\mathcal{M}}}\exp\left(2t\left(\Phi_{Y}\left(\frac{2\pi y}{q}\right)-1\right)\right)\leq\exp\left(\frac{6+12e^{4/e}+12e^{8/e}+6e^{32/e}}{\alpha^{1/4}}\right).$ (70)
@@ 6.4 Mixing Time Upper Bound
Now we will conclude the mixing time upper bound for the random walks on $n\times n$ contingency tables. For any $\varepsilon\in(0,1)$, set
$\alpha:=\inf\left\{z\geq 1:4\varepsilon^{2}\geq\exp\left(\frac{6+12e^{4/e}+12e^{8/e}+6e^{32/e}}{z^{1/4}}\right)-1\right\}.$
By the $\ell^{2}$-bound, (69), and (70), we conclude that when $t=\dfrac{n^{2}q^{2}\log(\sqrt{\alpha}(n-1))}{8\pi^{2}}\left(1-\dfrac{4}{3q}\right)^{-1}$,
@@ 6.5 Mixing Time Lower Bound
We will now compute the mixing time lower bound for the $n\times n$ contingency table walk in continuous time. We showed that the discrete-time increment $Y$ has equivariant coordinates, with each coordinate having a marginal variance of $\dfrac{4}{n^{2}}$. We also note that the off-diagonal entries of the correlation matrix take values of either $\dfrac{-1}{n-1}$ or $\dfrac{1}{(n-1)^{2}}$. Since at each jump the $n\times n$ contingency table random walk updates at most four coordinates, each by ±$\pm 1$, we have that $\|Y\|_{1}\leq 4$. Therefore, applying (13) and setting $r=4$, $\sigma^{2}=\dfrac{4(n-1)^{2}}{n^{2}}$, we obtain
$t_{\operatorname{mix}}(\varepsilon)\geq\frac{n^{2}q^{2}}{8\pi^{2}}\left(\log(n)+\log\left(1-\frac{1}{n}\right)+\frac{1}{2}\log\left(\frac{\varepsilon^{-1}-1}{5}\right)\right)$

### v2

@@ 6.3 Bounding Decay in Characteristic Function for Small $\theta$
In this subsection, we bound the decay of the characteristic function $\Phi_{Y}(\theta)$ in the regime $\|\theta\|_{\infty}\leq\dfrac{1}{\sqrt{q}}$.
@@ 6.3.1 The Schur Complement of a Submatrix of $\Psi$
For convenience, we write $A:=\dfrac{mn}{(m-1)(n-1)}\Psi$ and thus
@@
Fix any $\ell\in[m-1],k\in[n-1]$. Let $A^{*}=A^{*}(\ell,k)$ be the principal submatrix of $A$ that only contains indices from the first $\ell$ rows and $k$ columns of the contingency table. We further partition $A^{*}$ as
@@
Let $A^{*}=A^{*}(\ell,k)$ be as above. Then the Schur complement is
@@
[…]
@@
Figure 1: Diagram for an index $(i,k)$ with a shared $k$th column with $(\ell,k)$ in the submatrix containing the first $\ell$ rows and $k$ columns of the $m\times n$ contingency table. Note that this diagram is only for counting purposes, and the rows and columns may have been permuted.
[…]
@@
$(A^{*}_{11}v)_{(i,j)}$ $=-\frac{4}{\ell k}+\frac{2}{\ell}+\frac{2}{k}+\frac{\ell-2}{\ell}+\frac{k-2}{k}-\frac{2(\ell-2)}{\ell k}-\frac{2(k-2)}{\ell k}-\frac{(\ell-2)(k-2)}{\ell k}$
$=1.$
$(\ell,k)$(i,j)(i,j)k2k-$2\ell$-2--
Figure 2: Diagram for an index $(i,j)$ that does not share a row or a column with $(\ell,k)$ in the submatrix containing the first $\ell$ rows and $k$ columns of the $m\times n$ contingency table. Note that this diagram is only for counting purposes, and the rows and columns may have been permuted.
Therefore, we have $A^{*}_{11}v=A^{*}_{12}$ and thus $(A^{*}_{11})^{-1}A^{*}_{12}=v$. Now, since there are k1k-1 indices with row $\ell$ other than $(\ell,k)$, and $\ell-1$ indices with column $k$ other than $(\ell,k)$, a direct computation yields
@@ 6.3.2 The Correlation Condition
We now show that the correlation matrices $\Gamma_{n}$ for the random walk on the $m\times n$ contingency table meet the Correlation Condition.
@@
Let $(\Gamma_{n})_{n=1}^{\infty}$ be the sequence of correlation matrices corresponding to the contingency table walk on an $n\times n$ table. Then $\Gamma_{n}$ meets the Correlation Condition with the bounding function
$g(\alpha)=\frac{\kappa_{2}^{2}(1+2e^{4/e}+2e^{8/e}+e^{32/e})}{\kappa_{1}\alpha^{1/4}}.$
@@
$\Psi_{n}((i,j),(i^{\prime},j^{\prime}))=\begin{cases}\frac{4(m-1)(n-1)}{mn}&(i,j)=(i^{\prime},j^{\prime})\\ \frac{2(m-1)(n-1)}{mn}&i=i^{\prime}\text{ or }j=j^{\prime}\ \text{but not both}\\ \frac{(m-1)(n-1)}{mn}&i\neq i^{\prime}\text{ and }j\neq j^{\prime}\end{cases}.$
It follows that $\sup_{n}\|\Psi_{n}\|_{\infty}=4$ which confirms condition (i). In order to check condition (ii), we first specify an ordering of the elements of $\Psi_{n}$ in order to compute each respective Schur complement: the first n1n-1 elements will consist of the first row of the table (incrementing by column index), second n1n-1 elements will consist of row two in the same order, continuing for each subsequent row.
@@
so that $M_{n,(\ell,k)}^{(22)}\in{\mathbb{R}}$ corresponds to the element $(\ell,k)$. Then, since $M_{n,(\ell,k)}$ is a principal submatrix of $\Psi_{n,(\ell,k)}$, by a similar decomposition as (39),
@@
$M_{n,(\ell,k)}^{(22)}-M_{n,(\ell,k)}^{(21)}\left(M_{n,(\ell,k)}^{(11)}\right)^{-1}M_{n,(\ell,k)}^{(12)}=\left(\frac{(m-1)(n-1)}{mn}\right)\left(1+\frac{\ell+k+1}{\ell k}\right).$
@@
Now, if $\ell=1$ or $k=1$, say $\ell=1$, then note that $\Psi_{n,(1,k)}=\dfrac{2(m-1)(n-1)}{mn}(I_{k}+J_{k})$. So by (43),
$a_{\ell,k}^{(n)}=\left(\Psi_{n,(\ell,k)}^{(22)}-\Psi_{n,(\ell,k)}^{(21)}\left(\Psi_{n,(\ell,k)}^{(11)}\right)^{-1}\Psi_{n,(\ell,k)}^{(12)}\right)^{-1}\geq\frac{\ell k}{(\ell+1)(k+1)}$
@@
$\sum_{\ell=1}^{m-1}\sum_{k=1}^{n-1}\left(\frac{1}{\alpha(m-1)(n-1)}\right)^{a_{\ell,k}^{(n)}}$ $\leq\sum_{\ell=1}^{m-1}\sum_{k=1}^{n-1}\left(\frac{1}{\alpha(m-1)(n-1)}\right)^{\frac{\ell k}{(\ell+1)(k+1)}}$
$\leq\frac{1}{\alpha^{1/4}}\sum_{\ell=1}^{m-1}\sum_{k=1}^{n-1}\left(\frac{1}{mn}\right)^{\frac{\ell k}{(\ell+1)(k+1)}}$
$\leq\frac{\kappa_{2}^{2}(1+2e^{4/e}+2e^{8/e}+e^{32/e})}{\kappa_{1}\alpha^{1/4}}$
@@
$t=\dfrac{mnq^{2}\log(\alpha(m-1)(n-1))}{16\pi^{2}}\left(1-\dfrac{4}{3q}\right)^{-1}.$
@@
$\sum_{y\in{\mathcal{M}}}\exp\left(2t\left(\Phi_{Y}\left(\frac{2\pi y}{q}\right)-1\right)\right)\leq\exp\left(\frac{\kappa_{2}^{2}(6+12e^{4/e}+12e^{8/e}+6e^{32/e})}{\kappa_{1}\alpha^{1/4}}\right).$ (76)
@@ 6.4 Mixing Time Upper Bound
We will conclude the mixing time upper bound for the random walks on $m\times n$ contingency tables. For any $\varepsilon\in(0,1)$, set
$\alpha:=\inf\left\{z\geq 1:4\varepsilon^{2}\geq\exp\left(\frac{\kappa_{2}^{2}(6+12e^{4/e}+12e^{8/e}+6e^{32/e})}{\kappa_{1}z^{1/4}}\right)-1\right\}.$
By the $\ell^{2}$-bound, (75), and (76), when $t=\dfrac{mnq^{2}\log(\alpha(m-1)(n-1))}{16\pi^{2}}\left(1-\dfrac{4}{3q}\right)^{-1}$,
@@ 6.5 Mixing Time Lower Bound
We will now compute the mixing time lower bound for the $m\times n$ contingency table walk in continuous time. We showed that the discrete-time increment $Y$ has equivariant coordinates, with each coordinate having a marginal variance of $\dfrac{4}{mn}$. We also note that the off-diagonal entries of the correlation matrix take values of $\dfrac{-1}{m-1},\dfrac{-1}{n-1}$, or $\dfrac{1}{(m-1)(n-1)}$. Since at each jump the $m\times n$ contingency table random walk updates at most four coordinates, each by $\pm 1$, we have that $\|Y\|_{1}\leq 4$. Therefore, applying (17) using that $\kappa_{1}n\leq m\leq\kappa_{2}n$, and setting $r=4$, $\sigma^{2}=\dfrac{4(m-1)(n-1)}{mn}$, we obtain
$t_{\operatorname{mix}}(\varepsilon)\geq\frac{mnq^{2}}{16\pi^{2}}\left(\log\left((m-1)(n-1)\right)+\log\left(\frac{\varepsilon^{-1}-1}{5}\right)\right)$
@@
7 Acknowledgments
The authors thank David Sivakoff for valuable discussions, feedback, and review of the manuscript.
ZF is supported by the Department of Mathematics of The Ohio State University, and AH is supported by the Department of Statistics of The Ohio State University, both via graduate teaching associateship.